# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ 4.2 Density Estimation from Data
We provide a brief summary of the framework for such adaptive transport maps in Appendix A. Here we note that among the several parameters defining the adaptive transport maps is the number of terms in the basis function expansion for each map component. Figure 5 illustrates the effect of varying this allowed maximum number of terms and shows that, for each of the class 0 and class 1 data, increasingly good approximations to the density can be obtained. Classification results obtained via the transport map approximated density distribution functions are also shown. For comparison, classification results obtained using several Matlab classifiers [25] are included in Figure 5 as well.
@@ 5 Application to Gene Expression Data
To conduct an initial assessment of the potential of measure transport techniques for phenotypic classification within the radiation biology domain, we use the GSE43151 data set [30] available from the Gene Expression Omnibus (GEO) database [9], as in previous work [22, 23]. This data set contains gene expression data for 121 human blood samples, some with zero exposure to radiation, and others exposed to low-dose radiation or high-dose radiation. Out of the total number of samples, 18 are in the zero-dose category, 87 are in the low-dose category, and 16 are in the high-dose category. For more details on the GSE43151 data set, the reader may refer to [30]. The data set, as used in our experiments and the previous works [22, 23], is available from [21].
@@
Genes correspond to the random variables jointly distributed according to the unknown target probability densities we seek to estimate. For feature (i.e., random variables) selection, we consider genes in biological pathways relevant to the GSE43151 data set, as determined by the pathway ranking procedure from [22, 23]. As previously mentioned in Section 4.1, we investigate two different measure transport techniques, namely, explicitly enforcing sparsity patterns in the transport maps and adaptive transport maps. Details appear in Section 5.1 for the former approach and in Section 5.2 for the latter one. In each case, the resulting transport maps are sparse, as defined towards the end of Section 4.1 and depicted in Figure 3.
@@
As discussed next in Section 5.1 below, prior knowledge extracted from known biological pathways, such as relationships among genes, could be used to fix sets of active variables (i.e., sets of variables that the map components depend on) when computing the sparse transport maps. This is advantageous computationally, particularly if the extracted relationships are known with (reasonable) certainty. However, this approach could end up being biased and too restrictive (for example, missing possible dependencies between genes that are a priori unknown as well as obscuring their change when exposed to radiation). Hence, in Section 5.2, we also consider adaptive transport maps without any a priori imposed restrictions on the sparsity patterns of their components. With the latter class of maps, via specifically designed algorithms [2], sparse transport maps can be constructed and computed without the need to fix the set of active variables a priori. For our application, adaptive transport maps may thus offer more flexibility and would be more advantageous when prior knowledge about gene relationships is lacking or uncertain.
@@ 5.1 Prescribed Sparsity Patterns from Gene Relationships
One simple way to incorporate prior information into transport map construction is by building a sparse transport map using knowledge of the dependence structure of the set of random variables under consideration. If the random variables adhere to a Bayesian network, then the active variables of the transport map may be obtained from this network in a straightforward way [40]. In doing so, the number of parameters that need to be fit can be significantly reduced compared with triangular and dense maps.
For the gene expression data under radiation exposure, we used the Kyoto Encyclopedia of Genes and Genome (KEGG) database [16] to extract Bayesian networks describing the relationships between the genes for a given pathway. While KEGG provides a variety of relationships between genes, we focused on protein-protein (PPrel) activation and inhibition relationships. In particular, if there was a directed edge from gene A to gene B in the KEGG graph of type PPrel-activation/inhibition, then an edge from A to B was added to the Bayesian network. We also added edges between genes that are part of a gene complex. This information was extracted from KEGG Markup Language (KGML) files and read into Python using the BioPython package. We only consider pathways that resulted in directed acyclic graphs (DAGs) under these PPrel relationships. Software to extract this information is available from [15]. The transport maps for the experiments reported in this section were computed using the TransportMaps library [28], version 2.0b3.
@@
To test out the performance of the sparse transport maps framework within this context, we selected 5 gene pathways from the KEGG database that performed well on the pathway activation scoring performed in [23] and resulted in DAGs as well, through the above analysis. We considered the following pathways: hsa05131, hsa04714, hsa04658, hsa05166, and hsa04650. We performed the classification task of discriminating zero-dose versus low-dose samples using the subset of genes corresponding to each pathway. We were interested in the small-data regime where prior information can provide the biggest benefit, so we used only 25% of the data for the training set and the rest for evaluation, resulting in 25 training data samples and 80 test samples. Since there is so little data in the zero-dose case (only 4 training samples) it makes sense to use a diagonal transport map for this case, while other approaches performed poorly. In the low dose-case, there are 21 training samples and we consider using the following maps:
@@
Refer to Figure 3 for a pictorial representation of the above listed types of maps. In each of the above cases, we have $T^{-1}\!:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, where $m$ is the number of genes extracted from the pathway.
Figure 6 displays the $F_{1}$ score [35] averaged over both classes for each of the pathways and each method including Naive Bayes as a baseline. From the plot, we see that transport maps with the sparsity pattern derived from prior knowledge provides an improvement in performance across all pathways except one. Note the poor performance of the dense triangular map, which is unable to fit such a small amount of data.
@@ 5.2 Adaptive Transport Maps Approach
We now move on to adaptive transport maps [2], for which transport map sparsity patterns (i.e., the dependence structure of the set of random variables jointly distributed according to the target probability density) are not being prescribed a priori. First, we continue with classification experiments in Section 5.2.1 below. Afterwards, in Section 5.2.2 we discuss a different application which shows the potential of adaptive transport maps to serve as a tool for scientific discovery.
@@ 5.2.1 Further Classification Experiments
To design classification experiments using probability distribution densities approximated via adaptive transport maps (as opposed to Section 5.1 above, where transport maps sparsity patterns were prescribed a priori), we selected two pathways from the KEGG database [16], namely pathways hsa04120 and hsa05202. This was based on the pathway ranking procedure from [23], using the GSE43151 data set, where pathway hsa04120 ranked as the top pathway showing significant differential activation in response to low-dose radiation, as compared to zero-dose, and pathway hsa05202 ranked as the top pathway showing significant differential activation in response to high-dose radiation, as compared to zero-dose. After accounting for genes present in both the GSE43151 data set and in the selected pathways, we were left with 108 genes for pathway hsa04120 and 98 genes for pathway hsa05202.
Within the context of the classification problem from Section 2, we performed two groups of experiments. The first was for classifying samples between zero-dose (class 0) and low-dose (class 1) radiation, using the set of 108 genes common to both the GSE43151 data set and the hsa04120 pathway to define the set of random variables for the class-conditional densities. In the second group of experiments, classification was between zero-dose (class 0) and high-dose (class 1) radiation, using the set of 98 genes common to both the GSE43151 data set and the hsa05202 pathway as the set of random variables. For each of these two “class 0 - class 1 scenarios” we then decided on a number of maximum terms allowed in the basis function expansion for each map component (see the end of Section 4.2) and, given a maximum number of terms allowed (more on our choices later), a sequence of runs, say $N$ of them, was carried out. For each of these runs:
@@
    An adaptive transport map approximation to each class-conditional density was computed.
@@
    Classification was done, per (1), using the resulting class-conditional densities and the samples held for inference.
@@
This is formally described in Algorithm 1, using the mathematical notation introduced previously in Section 3.
@@
    $\pi_{k}$: (unknown) target probability density for class $k\in\{0,\ldots,K-1\}$
    $Y_{k}=\{y\in\mathbb{R}^{m}\,:\,y\sim{\nu_{\pi_{k}}}\}$: set of all samples for class $k\in\{0,1,\ldots,K-1\}$
@@
7:         Compute an adaptive transport map approximation $T_{k}$ from the
         reference $\rho$ to the target density ${\pi_{k}}$ using only $\sigma\left(Y_{k}\right)^{\mathrm{tr}}$ as a training set.
@@
          $\nu_{\pi_{k}}\left(\sigma\left(Y_{k}\right)^{\mathrm{tr}}\right)=\nu_{\rho}({T_{k}^{-1}}\left(\sigma\left(Y_{k}\right)^{\mathrm{tr}}\right))$. (cf. equation (2))
@@
10:         Compute $\widehat{\pi}_{k}=\lfloor\frac{n_{k}}{2}\rfloor\diagup\sum_{s=0}^{K-1}\lfloor\frac{n_{s}}{2}\rfloor$, the estimates for the class prior
@@
14:              Compute the probability $\mathrm{Pr}(C=k\,|\,\mathcal{Y}=y)={\widehat{\pi}_{k}\,\pi_{k}(y)}\diagup{\sum_{s=0}^{K-1}\widehat{\pi}_{s}\,\pi_{s}(y)}$
              that sample $y$ belongs to class $k$ (cf. equation (1)), where $\pi_{s}(y)$ is
@@
17:     (Optional) For comparison, compute classification with Matlab’s naive Bayes
@@
     data sets from line 6 (Algorithm 1).
@@
Figure 7: Summary of classification results from adaptive transport maps (ATM) approximations to the class-conditional densities, naive Bayes (NB) classifier, and support vector machine (SVM) classifier. Zero-dose radiation is class 0; Low-dose radiation is class 1. The available number of samples were 18 for class 0 and 87 for class 1. For each experiment (i.e., run), the available samples were divided randomly into 50% for training and 50% for inference. The total number of experiments was 100. For this data set, all methods performed well, with SVM achieving perfect results in most experiments. On average, the classification results for the ATM density approximations and for naive Bayes were equal, as seen in Subfigure 7a. For disjoint subsets of experiments (Subfigures 7b-7d), the ATM and naive Bayes classifiers performed equally, or one slightly better than the other. For all experiments, classification results were more accurate for class 1. This is not surprising, since there were more data samples available for class 1.
@@
Results from the above described runs are summarized in the form of ${F}_{1}$ scores for the classification results. These are shown in Figure 7 for the case when class 0 and class 1 were set, respectively, as the zero-dose and low-dose radiation classes. For the zero-dose and high-dose classification scenario, results are summarized in Figure 8.
We note that for the experiments conducted, all classifiers ended up performing well. This was the case also in other preliminary experiments with the selected GSE43151 data set. A small number of terms (1-2 for the zero- and high-dose radiation classes; 1-6 for the low-radiation class) in the adaptive transport maps basis function expansions were sufficient for the approximations. Varying the maximum number of allowed terms in the transport maps basis function expansions did not yield appreciable differences in the results, contrary to the (non-biology) example from Figure 5. However, our experiments (as previous works from others) illustrate that the same measure transport framework can be used successfully (in practice) for diverse data sets.
@@ 5.2.2 Adaptive Transport Maps as a Tool for Scientific Discovery
As shown in the original publication [2], adaptive transport maps approximations can reveal (unknown) dependence structure among the set of random variables jointly distributed according to the target probability density. Thus, adaptive transport maps have the potential to serve as tools for scientific discovery. For our radiation biology application, a natural question would then be whether the sparsity patterns resulting from the transport maps approximations to the target densities contain or reveal any biological information, for example, relationships between the genes associated with the transport map sparsity patterns. This is another task within the framework introduced in Section 2 (see Figure 1), independent of the classification problem, where the measure transport approach can have an impact.
@@
For instance, the sparsity pattern shown in Figure 3b resulted from the computation of the transport maps needed for estimating the class-conditional density for low-dose radiation, using a subset of five genes from the GSE43151 data set as the random variables jointly distributed according to the target measure. From the sparsity pattern in Figure 3b, one question that arises naturally is whether within the radiation biology domain there are known relationships that correspond to the exhibited sparsity pattern, or whether there may exist unknown relationships, among the genes associated with random variables $\mathcal{{Y}}_{3}$ and $\mathcal{{Y}}_{4}$ or between the genes associated with the random variables $\mathcal{{Y}}_{1}$, $\mathcal{{Y}}_{2}$, $\mathcal{{Y}}_{3}$ and $\mathcal{{Y}}_{5}$. Thus there is potential for hypotheses generation and subsequent expert verification. In addition, within the context of the computational workflow from Figure 1, the resulting sparsity patterns may conceivably be useful for designing queries (regarding gene relationships) for large language models as well [33, 34].
@@
Now, given sufficient data samples, one can expect that sparsity patterns resulting from adaptive transport map density approximations to the data distribution will accurately uncover dependencies or associations hidden in the data. But even when data is scarce one may be able to use such sparsity patterns for hypotheses generation. Namely, one can gather statistics by computing many adaptive transport map density approximations using different (randomly selected) subsets of the data as training sets, and then identifying common dominating patterns appearing in the resulting collection of transport map sparsity patterns. We discuss this line of research next and in Figures 913, where results are shown from our computational experiments. In order to develop some intuition into the dependence structure of the genes from the GSE43151 data set, without assuming any knowledge that fixes the set of transport map active variables a priori, we consider a subset of six genes from the GSE43151 data set as the random variables jointly distributed according to the unknown target measure. That is, in what follows, $T^{-1}\!:\mathbb{R}^{6}\rightarrow\mathbb{R}^{6}$ in equation (17). For reference, information about the selected subset of genes is provided in Table 1 and Figure 14, both appearing in Appendix B.
@@
Figure 9: Due to the small number of total samples available, resulting transport map sparsity patterns may differ when different training subsets of the total samples set are used to compute adaptive transport maps. Subfigures 9a9c exemplify this for three select subsets (of 9 samples each) from the zero-dose radiation class. Similarly for the low-dose radiation class in Subfigures 9d9f, where each subset had 44 samples when computing the corresponding adaptive transport map. Statistics collected from various runs may nevertheless allow for uncovering information hidden in the data, as shown in Figures 1013.
@@
Recall from Figure 3 that in our pictorial representation of transport map sparsity patterns, the horizontal axis indexes the random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ and the vertical axis indexes the map components $\{(T^{-1})_{i}\}_{i=1}^{6}$. A square appears at the intersection of grid point $(j,i)$ if the $i$-th map component $(T^{-1})_{i}$ depends on the value $y_{j}$ of the $j$-th random variable $\mathcal{{Y}}_{j}$. For any given map component, the set of active variables is the set of random variables that the map component depends on. The active variables define a sparsity pattern for the transport map. We emphasize again that the adaptive transport map computation methods from [2] determine the set of active variables, without specification of it prior to computing the transport map.
@@
Due to the small number of total samples available in the GSE43151 data set, sparsity patterns resulting from adaptive transport map approximations to the target measure may differ when different subsets of the total samples set are used to compute the transport maps. Figure 9 exemplifies this for the zero-dose and low-dose radiation classes. However, we hypothesize that statistics collected from various runs may nevertheless allow for uncovering information hidden in the data. The computations we perform to collect such statistics are described in Algorithm 2.
@@
    $N$: number of runs (transport maps) in a series of computations of transport maps
    $T_{k}^{(N,r)}$: $r$-th adaptive transport map computed in the series of $N$ runs for Class $k$

### v4

@@ 4.2 Density Estimation from Data
We provide a summary of the framework for such adaptive transport maps in Appendix A. Here we note that among the several parameters defining the adaptive transport maps is the number of terms in the basis function expansions (47) used in the functional representation of each transport map component. Figure 5 illustrates the effect of varying this allowed maximum number of terms and shows that, for each of the class 0 and class 1 data, increasingly good approximations to the target densities can be obtained. Classification results obtained via the transport map approximated density distribution functions, using the classification formula (4), are also shown. For comparison, classification results obtained using a neural network classifier from Matlab and the naive Bayes and support vector machine Matlab classifiers [25] are included in Figure 5 as well.
@@ 5 Application to Gene Expression Data
As we discussed in Section 2, one of the problems motivating the present research is that of probability distribution density estimation for phenotypic classification problems, for example, for inferring class labels, such as radiation exposure levels, given sample data. Recall from Section 1.1 that we are concerned with data sets containing a small number of samples, specifically of order $O(1)$$O(10)$ samples, which are common in the radiation biology domain.
@@
To conduct an initial assessment of the potential of measure transport techniques for phenotypic classification within the above mentioned context, we use the GSE43151 data set [30] available from the Gene Expression Omnibus (GEO) database [9], as in previous work [22, 23]. This data set contains gene expression data for 121 human blood samples, some with zero exposure to radiation, and others exposed to low-dose radiation or high-dose radiation. Out of the total number of samples, 18 are in the zero-dose category, 87 are in the low-dose category, and 16 are in the high-dose category. For more details on the GSE43151 data set, the reader may refer to [30]. The data set, as used in our experiments and the previous works [22, 23], is available from [21].
@@
Genes correspond to the random variables jointly distributed according to the unknown target probability densities we seek to estimate. For feature (i.e., random variables) selection, we consider genes in biological pathways relevant to the GSE43151 data set, as determined by the pathway ranking procedure from [22, 23]. As previously mentioned at the end of Section 4.1, we investigate two different measure transport techniques, namely, explicitly enforcing sparsity patterns in the transport maps (i.e., explicitly enforcing a pre-determined dependence structure among the random variables in the target measure) and adaptive transport maps (for which a dependence structure among the set of random variables in the target measure will not be enforced a priori). Details appear in Section 5.1 for the former approach and in Section 5.2 for the latter one. In each case, the resulting transport maps are sparse, as defined towards the end of Section 4.1 and depicted in Figure 3. Recall from the discussion in Section 4.1, in particular, that sparsity of a triangular transport map is a property of the target probability measure and not a property of the number of samples used for learning the transport map.
@@
As discussed next in Section 5.1 below, prior knowledge extracted from known biological pathways, such as relationships among genes, could be used to fix sets of active variables (i.e., sets of variables that the map components depend on), and thereby explicitly enforce sparsity patterns in the transport maps when learning the maps. This is advantageous computationally, particularly if the extracted relationships among genes are known with (reasonable) certainty. However, this approach could end up being biased and too restrictive (for example, missing possible dependencies between genes that are a priori unknown as well as obscuring their change when exposed to radiation). Hence, in Section 5.2, we also consider adaptive transport maps without any a priori imposed restrictions on the sparsity patterns of the transport map components. With adaptive transport maps, via specifically designed algorithms [3], sparse transport maps can be constructed and learned without the need to fix the set of active variables a priori. (Refer to the discussion surrounding equation (25) and also to Appendix A.) For our application, adaptive transport maps may thus offer more flexibility and would be more advantageous when prior knowledge about gene relationships is lacking or uncertain.
@@ 5.1 Prescribed Sparsity Patterns from Gene Relationships
One simple way to incorporate prior information into transport map construction is by building a sparse transport map using knowledge of the dependence structure of the set of random variables from the target measure of interest. If the random variables adhere to a Bayesian network, then the active variables of the transport map (as defined in Section 4.1 – see Figure 3) may be obtained from this network in a straightforward way [39]. In doing so, the number of parameters that need to be fit in order to learn the transport map can be significantly reduced compared with triangular dense maps.
For the gene expression data under radiation exposure, we used the Kyoto Encyclopedia of Genes and Genome (KEGG) database [16] to extract Bayesian networks describing the relationships between the genes for a given pathway. While KEGG provides a variety of relationships between genes, we focused on protein-protein (PPrel) activation and inhibition relationships. In particular, if there was a directed edge from gene A to gene B in the KEGG graph of type PPrel-activation/inhibition, then an edge from A to B was added to the Bayesian network. We also added edges between genes that are part of a gene complex. This information was extracted from KEGG Markup Language (KGML) files and read into Python using the BioPython package. We only considered pathways that resulted in directed acyclic graphs (DAGs) under these PPrel relationships. Software to extract this information is available from [15]. The transport maps for the experiments reported in this section were computed using the TransportMaps library [28], version 2.0b3.
@@
To test out the performance of the sparse transport maps framework within this context, we selected 5 gene pathways from the KEGG database that performed well on the pathway activation scoring performed in [23] and resulted in DAGs as well, through the above analysis. We considered the following pathways: hsa05131, hsa04714, hsa04658, hsa05166, and hsa04650. We performed the classification task of discriminating zero-dose versus low-dose samples using the subset of genes from the GSE43151 data set corresponding to each pathway. We were interested in the “small-data” (i.e., small number of samples) regime where prior information can provide the biggest benefit, so we used only 25% of the total number of data samples for the training set and reserved the rest of the samples for the inference set. Since there is so little data in the zero-dose case (only 4 training samples) it made sense to use a diagonal transport map for this case, because other approaches performed poorly. In the low-dose case, there are 21 training samples and we consider using the following maps:
@@
Refer to Figure 3 from Section 4.1 for a pictorial representation of the above listed types of maps. In each of the above cases, we have $T^{-1}\!:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, where $m$ is the number of genes extracted from the pathway.
Figure 6 displays on the vertical axis the $F_{1}$ classification score [35]
$F_{1}$ $=$ $2\,\times\,\frac{\mathrm{precision}\times\mathrm{recall}}{\mathrm{precision}+\mathrm{recall}}\,,$ (26)
averaged over both classes for each of the pathways (note that pathways are indexed on the horizontal axis in Figure 6) and each transport map method considered, and including naive Bayes classification as a baseline. In equation (26), for each class $k\in\{0,1\}$ (0: zero-dose class; 1: low-dose class), $\mathrm{precision}=\mathrm{TP}/(\mathrm{TP+FP})$ and $\mathrm{recall}=\mathrm{TP}/(\mathrm{TP+FN})$, where $\mathrm{TP}$ (True Positive) is the total number of samples correctly classified as belonging to class $k$, $\mathrm{FP}$ (False Positive) is the total number of samples incorrectly classified as belonging to class $k$, and $\mathrm{FN}$ (False Negative) is the total number of samples incorrectly classified as not belonging to class $k$. From Figure 6, we see that transport maps with the sparsity pattern derived from prior knowledge provides an improvement in performance across all pathways except one. Note the poor performance of the dense triangular map, for which we were unable to fit adequately such a small number of data samples.
@@ 5.2 Adaptive Transport Maps Approach
We now move on to adaptive transport maps [3], for which transport map sparsity patterns (i.e., the dependence structure of the set of random variables jointly distributed according to the target measure) are not being prescribed a priori. First, we continue with classification experiments in Section 5.2.1 below. Afterwards, in Section 5.2.2 we discuss a different application which shows the potential of adaptive transport maps to serve as a tool for scientific discovery.
@@ 5.2.1 Further Classification Experiments
To design classification experiments using probability distribution densities approximated via adaptive transport maps (as opposed to Section 5.1 above, where transport maps sparsity patterns were prescribed a priori), we selected two pathways from the KEGG database [16], namely pathways hsa04120 and hsa05202. This pathway selection was based on the pathway ranking procedure from [23], using the GSE43151 data set, where pathway hsa04120 ranked as the top pathway showing significant differential activation in response to low-dose radiation, as compared to zero-dose, and pathway hsa05202 ranked as the top pathway showing significant differential activation in response to high-dose radiation, as compared to zero-dose. After accounting for genes present in both the GSE43151 data set and in the selected pathways, we were left with 108 genes for pathway hsa04120 and 98 genes for pathway hsa05202.
Within the context of the classification problem via formula (4) from Section 2, we performed two groups of experiments. The first was for classifying samples between zero-dose (class 0) and low-dose (class 1) radiation, using the set of 108 genes common to both the GSE43151 data set and the hsa04120 pathway to define the set of random variables. In the second group of experiments, classification was between zero-dose (class 0) and high-dose (class 1) radiation, using the set of 98 genes common to both the GSE43151 data set and the hsa05202 pathway as the set of random variables. For each of these two “class 0 - class 1 scenarios” we then decided on a number of maximum terms allowed in the basis function expansions (47) for each map component (refer to the discussion at the end of Section 4.2 and also to Figure 5) and, given a maximum number of terms allowed (more on our choices later), a sequence of runs, say $N$ of them, was carried out. For each of these runs:
@@
    An adaptive transport map approximation to each target class-conditional density was computed.
@@
    Classification was done, per the classification formula (4), using the resulting transport map approximations to the class-conditional target densities and the samples held for inference.
@@
This is formally described in Algorithm 1, using the mathematical notation introduced previously in Section 3. The outcomes from the above described runs are summarized in the form of ${F}_{1}$ scores (26) for the classification results. These are shown in Figure 7 for the case when class 0 and class 1 were set, respectively, as the zero-dose and low-dose radiation classes. For the zero-dose and high-dose classification scenario, results are summarized in Figure 8.
@@
    $\psi_{k}$: (unknown) target probability density for class $k\in\{0,\ldots,K-1\}$
    $Y_{k}=\{y\in\mathbb{R}^{m}\,:\,y\sim{\nu_{\psi_{k}}}\}$: set of all samples for class $k\in\{0,1,\ldots,K-1\}$
@@
7:         Learn an adaptive transport map approximation $T_{k}$ from the
         reference $\rho$ to the target density ${\psi_{k}}$ using only $\sigma\left(Y_{k}\right)^{\mathrm{tr}}$ as a training set.
@@
          $\nu_{\psi_{k}}\left(\sigma\left(Y_{k}\right)^{\mathrm{tr}}\right)=\nu_{\rho}({T_{k}^{-1}}\left(\sigma\left(Y_{k}\right)^{\mathrm{tr}}\right))$. (cf. equation (5))
@@
10:         Compute $\widehat{\psi}_{k}=\lfloor\frac{n_{k}}{2}\rfloor\diagup\sum_{s=0}^{K-1}\lfloor\frac{n_{s}}{2}\rfloor$, the estimates for the class prior
@@
14:              Compute the probability $\mathrm{Pr}(C=k\,|\,\mathcal{Y}=y)={\widehat{\psi}_{k}\,\psi_{k}(y)}\diagup{\sum_{s=0}^{K-1}\widehat{\psi}_{s}\,\psi_{s}(y)}$
              that sample $y$ belongs to class $k$ (cf. equation (4)), where $\psi_{s}(y)$ is
@@
17:     (Optional) For comparison, perform classification with Matlab’s naive Bayes
@@
     data sets from line 6 above.
@@
Figure 7: Summary of classification results from adaptive transport maps (ATM) approximations to the class-conditional densities, naive Bayes (NB) classifier, and support vector machine (SVM) classifier. Zero-dose radiation is class 0; Low-dose radiation is class 1. The available number of samples were 18 for class 0 and 87 for class 1. For each experiment (i.e., run), the available samples were divided randomly into 50% for training and 50% for inference. The total number of experiments was 100. For this data set, all methods performed well, with SVM achieving perfect results in most experiments. On average, the classification results for the ATM density approximations and for naive Bayes were equal, as seen in Subfigure 7a. For disjoint subsets of experiments (Subfigures 7b-7d), the ATM and naive Bayes classifiers performed equally, or one slightly better than the other. For all experiments, classification results were more accurate for class 1. This is not surprising since there were more data samples available for transport map training for class 1.
@@
We note that for the experiments conducted, all classifiers ended up performing well. This was the case also in other preliminary experiments with the selected GSE43151 data set. A small number of terms (1-2 for the zero- and high-dose radiation classes; 1-6 for the low-radiation class) in the adaptive transport maps basis function expansions (47) (refer to the discussion at the end of Section 4.2 and also to Figure 5) were sufficient for the approximations. Varying the maximum number of allowed terms in the transport maps basis function expansions (47) did not yield appreciable differences in the results, contrary to the (non-biology) example from Figure 5. However, our experiments (as previous works from others) illustrate that the same measure transport framework can be used successfully (in practice) for diverse data sets.
@@ 5.2.2 Adaptive Transport Maps as a Tool for Scientific Discovery
We now depart from the phenotypic classification problems of Sections 5.1 and 5.2.1 and consider the problem of inferring the dependence structure among random variables in the unknown target measure, which was previously mentioned in Section 2 as being also of interest to us.
@@
As shown in the original publication [3], adaptive transport maps approximations can reveal (unknown) dependence structure among the set of random variables jointly distributed according to the target probability density. Thus, adaptive transport maps have the potential to serve as tools for scientific discovery. Recall from Section 4.1 (refer also to Figure 3) that information about the dependence structure of the random variables in the target measure is encoded in the triangular transport map sparsity patterns learned from data samples. (That is, triangular transport map sparsity patterns reflect structure in the target space, and should not be thought of as related to a statement about the number of training samples used to learn the transport maps.) For our radiation biology application, a natural question would then be whether the sparsity patterns resulting from the transport maps approximations to the target densities contain or reveal any biological information, for example, relationships between the genes associated with the transport map sparsity patterns. This is another task within the workflow introduced in Section 2, independent of the classification problems already explored in Sections 5.1 and 5.2.1, where the measure transport approach can have an impact.
@@
For instance, the transport map sparsity pattern shown in Figure 3b resulted from learning a transport map for estimating the class-conditional density for low-dose radiation, trained using a subset of five genes from the GSE43151 data set as the random variables jointly distributed according to the target measure. From the sparsity pattern in Figure 3b, one question that arises naturally is whether within the radiation biology domain there are known relationships that correspond to the exhibited transport map sparsity pattern, or whether there may exist unknown relationships, among the genes associated with random variables $\mathcal{{Y}}_{3}$ and $\mathcal{{Y}}_{4}$ (see row for 4th map component in Figure 3b) or between the genes associated with the random variables $\mathcal{{Y}}_{1}$, $\mathcal{{Y}}_{2}$, $\mathcal{{Y}}_{3}$ and $\mathcal{{Y}}_{5}$ (see row for 5th map component in Figure 3b). Thus there is potential for hypotheses generation and subsequent expert verification. In addition, within the context of the computational workflow from Figure 2, the resulting transport map sparsity patterns may conceivably be useful for designing queries (regarding gene relationships) for large language models as well [33, 34].
Now, given a sufficient number of data samples, one can expect that a triangular transport map sparsity pattern resulting from an adaptive transport map density approximation to the data distribution will accurately uncover information hidden in the data about dependencies or associations among the random variables. However, when the number of training samples is limited in amount ($O(1)$$O(10)$ samples in our case), it may be difficult to infer dependence structure in the target measure by learning a single adaptive transport map, as we discuss in Figure 9. Nevertheless, even when the number of training samples is limited in amount, we hypothesize that one may be able to use adaptive transport map sparsity patterns for inferring dependency structure among random variables in the target measure.
@@
More precisely, we propose a randomized technique designed to gather statistics by computing a series of adaptive transport map density approximations using different (randomly selected) subsets of the sample data as training sets, and then identifying common dominating patterns appearing in the resulting collection of transport map sparsity patterns. We discuss this line of research next and in Figures 913, where results are shown from our computational experiments. In order to develop some intuition into the dependence structure of the genes from the GSE43151 data set, without assuming any knowledge that fixes the set of transport map active variables a priori, we consider a subset of six genes from the GSE43151 data set as the random variables jointly distributed according to the unknown target measure. That is, in what follows we have $T^{-1}\!:\mathbb{R}^{6}\rightarrow\mathbb{R}^{6}$ (analogously, $m=6$ in equation (22)). For reference, information about the selected subset of genes is provided in Table 1 and Figure 14, both appearing in Appendix B.
@@
Figure 9: Due to the small number of total samples available for training, resulting adaptive transport map sparsity patterns may differ when different training subsets of the total samples set are used to learn adaptive transport maps. Subfigures 9a9c exemplify this for three select subsets (of 9 training samples each) from the zero-dose radiation class. Similarly for the low-dose radiation class in Subfigures 9d9f, where each subset had 44 training samples for learning the corresponding adaptive transport map. Statistics collected from sparsity pattens from a series of learned adaptive transport maps may nevertheless allow for inferring from data information about the dependency structure among random variables in the target measure, as shown in Figures 1013.
@@
Recall from Figure 3 that in our pictorial representation of transport map sparsity patterns, the horizontal axis indexes the random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{6}$ and the vertical axis indexes the map components $\{(T^{-1})_{i}\}_{i=1}^{6}$. A square appears at the intersection of grid point $(j,i)$ if the $i$-th map component $(T^{-1})_{i}$ depends on the value $y_{j}$ of the $j$-th random variable $\mathcal{{Y}}_{j}$. For any given map component, the set of active variables is the set of random variables that the map component depends on. The active variables define a sparsity pattern for the transport map. We emphasize again that the adaptive transport map computation methods from [3] (summarized in Appendix A of the present paper) determine the set of active variables, without specification of it prior to learning the transport map.
@@
Due to the small number of total samples available in the GSE43151 data set (18 in the zero-dose category, 87 in the low-dose category, and 16 in the high-dose category – see the beginning of Section 5), transport map sparsity patterns resulting from adaptive transport map approximations to the target measure may differ when different subsets of the total samples set are used to learn transport maps. Figure 9 exemplifies this for the zero-dose and low-dose radiation classes. However, as already noted, we hypothesize that statistics collected from sequences of learned adaptive transport maps may nevertheless allow for adaptive transport map sparsity patterns to be used in order to infer structure in the target measure from information hidden (or encoded) in the data. The computations we perform to collect such statistics are described in Algorithm 2 and outlined next.
@@
    $N$: number of transport maps in the series of adaptive transport maps to be learned
        (see line 3 below)
    $T_{k}^{(N,r)}$: $r$-th map in the series of $N$ adaptive transport maps learned for Class $k$